If a company has 20,000 shares of 5% preferred stock outstanding, selling for $72 per share, what is the cost of preferred stock?

A- 5.40%

B- 6.20%

C- 7.18%

D- 4.73%

E- 6.94%

Please show the calculation steps so I can understand. Thank you!
Answer #1

Assuming face value to be $100

Annual dividend = 0.05 * 100 = 5

Cost of preferred stock = (Annual dividend / share price) * 100

Cost of preferred stock = (5 / 72) * 100

Cost of preferred stock = 6.94%